2(x+8)-7=2+4(X-4) whats the answer
Answer by checkley77(12844)

2(x+8)-7=2+4(X-4)
2x+16-7=2+4x-16
2x-4x=2-16-16+7
-2x=-32+9
-2x-23
x=23/-2
x=-11.5 ans.
Proof:
2(11.5+8)-7=2+4(11.5-4)
2*19.5-7=2+4*7.5
39-7=2+30
32=32